OpenStudy (anonymous):

10m^4/9a^2 times 3a^2b

OpenStudy (anonymous):

is it 3*a^(2b) or 3*a^2 * b?

OpenStudy (anonymous):

it's 3a^2b

OpenStudy (anonymous):

Is B part of a's exponent?

OpenStudy (anonymous):

3a to the second power times b

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok. So simplifying the coefficients separately you have: 10/9 * 3 = 30/9 = 10/3 m^4 remains. a^2/a^2 = 1 b remains, leaving: 10/3 * b*m^4

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

you're welcome